# Algorithms and Lower Bounds for the Maximum Overlap of Two Polygons Under Translation

### v1

@@ Abstract
A fundamental problem in shape matching and geometric similarity is computing the maximum area overlap between two polygons under translation. For general simple polygons, the best-known algorithm runs in $\mathcal{O}((nm)^{2}\log(nm))$ time [Mount, Silverman, Wu ’96], where $n$ and $m$ are the complexities of the input polygons. In a recent breakthrough that received the SoCG Best Paper Award 2025, Chan and Hair gave a linear-time algorithm for the special case when both polygons are convex. A key challenge in computational geometry is to design improved algorithms for other natural classes of polygons. We address this by presenting an $\mathcal{O}((nm)^{3/2}\log(nm))$-time algorithm for the case when both polygons are orthogonal, probably the most popular class of polygons besides convex and simple ones. This is the first algorithm for polygon overlap on orthogonal polygons that is faster than the almost 30 years old algorithm for general simple polygons.
Complementing our algorithmic contribution, we provide $k$-SUM lower bounds for problems on simple polygons with only orthogonal and diagonal edges. First, we establish that there is no algorithm for polygon overlap with running time $\mathcal{O}(\max(n^{2},nm^{2})^{1-\varepsilon})$, where $m\leq n$, unless the $k$-SUM hypothesis fails. This matches the running time of our algorithm when $n=m$. We use part of the above construction to also show a lower bound for the polygon containment problem, a popular special case of the overlap problem. Concretely, there is no algorithm for polygon containment with running time $\mathcal{O}(n^{2-\varepsilon})$ under the $3$-SUM hypothesis, even when the polygon to be contained has $m=\mathcal{O}(1)$ vertices. Our lower bound shows that polygon containment for these types of polygons (i.e., with diagonal edges) is strictly harder than for orthogonal polygons, and also strengthens the previously known $3$-SUM lower bound for polygon containment of [Barequet, Har-Peled ’01]. Furthermore, our lower bounds show tightness (up to polylogarithmic factors) of the algorithms of [Avnaim, Boissonnat ’89] and [Mount, Silverman, Wu ’96] when $m\in\mathcal{O}(1)$.
@@ 1 Introduction
We consider the problem of computing the maximum overlap of two polygons $P$ and $Q$ with $n$ and $m$ vertices, respectively, i.e., finding a translation $\tau=(\tau_{x},\tau_{y})\in\mathbb{R}^{2}$ such that the area of $P\cap(Q+\tau)$ is maximum. This can be seen as a formalization of matching the shapes $P$ and $Q$, and the maximum overlap is a natural similarity measure: the larger it is relative to the areas of $P$ and $Q$, the more similar they are. The problem was first studied by Mount, Silverman and Wu [16] for general simple polygons. They show that the overlap function $\tau\mapsto{{\color[rgb]{0,0,0}\mathrm{area}}}(P\cap(Q+\tau))$ is a continuous and piecewise polynomial surface of degree at most two, consisting of ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{2})$ pieces. Furthermore, they give an example of two orthogonal polygons (for arbitrary values of $n$ and $m$) that realize the bound, i.e., the complexity of the overlap function is $\Omega((nm)^{2})$, and hence the bound is tight in the worst-case. They also describe an algorithm to compute the entire function within the same bound ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{2})$ on the running time, which yields an ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{2}\log(nm))$ algorithm for the polygon overlap problem. They posed the open question whether an algorithm with better running time exists. Approximation algorithms for finding the maximum overlap are described in [8, 7, 13]. See also the survey [17] on polygons from the handbook of discrete and computational geometry for an overview.
@@
Convex Containment $\Omega(n+m)$ ${{\color[rgb]{0,0,0}\mathcal{O}}}(n+m)$ [5]
MaxOverlap $\Omega(n+m)$ ${{\color[rgb]{0,0,0}\mathcal{O}}}(n+m)$ [5]
Orthogonal Containment ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{1-{{\color[rgb]{0,0,0}\mathrm{\varepsilon}}}})$ [15] ${{\color[rgb]{0,0,0}\tilde{\mathcal{O}}}}(nm)$ [14]
MaxOverlap ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{1-{{\color[rgb]{0,0,0}\mathrm{\varepsilon}}}})$ [15] ${{\color[rgb]{0,0,0}\tilde{\mathcal{O}}}}((nm)^{3/2})$
@@
Containment ${{\color[rgb]{0,0,0}\mathcal{O}}}(n^{2-{{\color[rgb]{0,0,0}\mathrm{\varepsilon}}}})$ ${{\color[rgb]{0,0,0}\tilde{\mathcal{O}}}}((nm)^{2})$ [2]
@@
MaxOverlap ${{\color[rgb]{0,0,0}\mathcal{O}}}(\max(n^{2},nm^{2})^{1-{{\color[rgb]{0,0,0}\mathrm{\varepsilon}}}})$ ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{2})$ [2]
@@
In this paper, we consider the problem for another well-studied subclass of simple polygons, where both $P$ and $Q$ are orthogonal polygons, i.e., simple polygons where every edge is either horizontal or vertical:
@@
For this problem, we give the first improvement over the nearly three decades old ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{2}\log(nm))$ bound from [16]. In particular, we prove the following theorem.
@@
The corresponding hardness assumption in fine-grained complexity theory is the $k$-SUM-Hypothesis [1, 12]:
@@
Our reduction can also be modified to obtain a lower bound for the polygon containment problem. While this problem is a special case of the maximum overlap problem, it has attracted more attention than the overlap problem. In the containment problem we want to know whether there is some $\tau\in\mathbb{R}^{2}$ so that $(Q+\tau)\subset P$ and thus ${{\color[rgb]{0,0,0}\mathrm{area}}}(P\cap(Q+\tau))={{\color[rgb]{0,0,0}\mathrm{area}}}(Q)$. This problem was first studied by Chazelle in 1983 [6], followed by numerous other results. Of particular interest, we only mention the results when both $P$ and $Q$ are simple or orthogonal polygon and only translations are allowed; a plethora of results exist when rotations are allowed and for more restricted variants such as when one or both of the polygons are convex; for exhaustive literature reviews on the problem, see the recent papers [15, 4]. Avnaim and Boissonnat [2] gave an algorithm for solving polygon containment of two simple polygons in ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{2}\log(nm))$ time. Hernández-Barrera [14] improved the running time to ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{2})$ and solved the case where both $P$ and $Q$ are orthogonal in ${{\color[rgb]{0,0,0}\mathcal{O}}}(nm\log(nm))$ time.
As for known lower bounds, Barequet and Har-Peled [3] showed that when $P$ and $Q$ both have $n$ vertices, no algorithm has running time ${{\color[rgb]{0,0,0}\mathcal{O}}}(n^{2-\varepsilon})$ under the $3$-SUM hypothesis. Künnemann and Nusser [15] gave various conditional lower bounds for polygon containment and the more general largest scaled containment problem, which is to find the largest homothet (i.e., scaled copy) of $Q$ that fits in $P$. For polygon containment, they showed that when $P$ and $Q$ are orthogonal and have arbitrary numbers $n$ and $m$ of vertices, then no algorithm has running time ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{1-\varepsilon})$ time under the Orthogonal Vectors (OV) hypothesis. Assuming the $4$-SUM hypothesis, they showed that for the largest scaled containment problem, no algorithm has running time ${{\color[rgb]{0,0,0}\mathcal{O}}}(n^{2-\varepsilon})$, even when both $P$ and $Q$ are orthogonal and $m=O(1)$.
@@
  • The arrangement that is constructed for the polygon overlap problem in [16] has worst-case size $\Theta((nm)^{2})$, even for orthogonal polygons. We break this natural barrier with our ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{3/2}\log nm)$ running time.
@@
  • With our improved lower bounds, we show that the decades-old algorithms of [2, 16, 14] for polygon overlap and containment are all tight up to polylogarithmic factors for $m\in{{\color[rgb]{0,0,0}\mathcal{O}}}(1)$.
@@ 1.1 Technical Overview
Let $P$ and $Q$ be two orthogonal polygons with $n$ and $m$ vertices respectively. We give the first algorithm improving on the original ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{2})$ algorithm from Mount, Silverman and Wu [16] for non-convex polygons. That is, for the subclass of orthogonal polygons we give an ${{\color[rgb]{0,0,0}\tilde{\mathcal{O}}}}((nm)^{3/2})$ algorithm.
@@
Similar to the approach presented in [16], we consider the piecewise polynomial surface of degree at most two describing the function $\tau\mapsto{{\color[rgb]{0,0,0}\mathrm{area}}}(P\cap(Q+\tau))$. We observe that the function splits into ${{\color[rgb]{0,0,0}\mathcal{O}}}(nm)$ rectangles, each with their own bilinear function, and $\tau\mapsto{{\color[rgb]{0,0,0}\mathrm{area}}}(P\cap(Q+\tau))$ can be expressed as the sum of bilinear functions of rectangles containing $\tau$. We may now sweep over these rectangles to compute the translation maximizing the intersection of $P$ and $Q+\tau$ in ${{\color[rgb]{0,0,0}\tilde{\mathcal{O}}}}((nm)^{2})$ time. We improve this sweep-line algorithm with a batching technique which groups events for the sweep-line algorithm into ${{\color[rgb]{0,0,0}\mathcal{O}}}(\sqrt{nm})$ groups of size ${{\color[rgb]{0,0,0}\mathcal{O}}}(\sqrt{nm})$ in both $x$- and $y$-direction resulting in ${{\color[rgb]{0,0,0}\mathcal{O}}}(nm)$ small windows in which we identify the maximizing translation in ${{\color[rgb]{0,0,0}\tilde{\mathcal{O}}}}(\sqrt{nm})$ time as follows: Via algebraic manipulations we rewrite the partial sums of bilinear functions, as they would have been encountered and evaluated by the sweep-line algorithm, into a set $V$ of ${{\color[rgb]{0,0,0}\mathcal{O}}}(\sqrt{nm})$ vectors in ${{\color[rgb]{0,0,0}\mathbb{R}}}^{3}$. Identifying the maximum among the ${{\color[rgb]{0,0,0}\mathcal{O}}}(nm)$ candidate translations in the small window then reduces to ${{\color[rgb]{0,0,0}\mathcal{O}}}(\sqrt{nm})$ extreme-point queries on the set $V$ which, as $V\subset{{\color[rgb]{0,0,0}\mathbb{R}}}^{3}$, can be answered in ${{\color[rgb]{0,0,0}\tilde{\mathcal{O}}}}(\sqrt{nm})$ total time, via the convex hull of $V$. The result is an algorithm with running time in ${{\color[rgb]{0,0,0}\tilde{\mathcal{O}}}}((nm)^{3/2})$.
@@
Comparing our techniques to the algorithm given by Mount, Silverman and Wu [16], we observe that our approach is similar in spirit in that we consider the piecewise polynomial surface of degree at most two describing the function $\tau\mapsto{{\color[rgb]{0,0,0}\mathrm{area}}}(P\cap(Q+\tau))$. As noted in Section 1, Mount, Silverman and Wu showed that the overlap function can have complexity $\Omega((nm)^{2})$ even for orthogonal polygons, so their method of explicitly computing the polynomial surface is unlikely to yield improved bounds. Unlike their algorithm, however, we manage to circumvent the explicit computation of the polynomial surface represented as the sum of bilinear polynomials defined on rectangular regions via our novel recast to extreme-point queries to a three-dimensional convex hull. With this, we improve the running time of the algorithm given by Mount, Silverman and Wu by a factor of $\sqrt{nm}$.
@@
Comparing our techniques with the techniques of the 3-SUM hypothesis based lower bound from Barequet and Har-Peled [3] and the Orthogonal Vectors hypothesis based lower bound from Künnemann and Nusser [15] for the OrthMaxOverlap problem, we observe the following: Both their techniques essentially consider only one pair of combs, restricting the set $\tau=(\tau_{x},\tau_{y})$ of translations such that $Q+\tau$ is contained in $P$ to one of $nm$ possible $x$-translations. Combining this with a second set of combs restricting the $y$-translations is straight-forward, however, there is no interdependence between $\tau_{x}$ and $\tau_{y}$, as both combs can be aligned independently of one another, giving ${{\color[rgb]{0,0,0}\mathcal{O}}}(nm)$ instead of the desired ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{2})$ translations that ‘need to be checked’. This problem is addressed in the 4-SUM lower bound of [15] by considering scalings additionally to translations, to determine the biggest copy of $Q$ in $P$. This additional degree of freedom allows to encode more complexity, however, at the cost of showing a lower bound for a more general problem (hence, a weaker bound). In our work, we address this issue by introducing diagonal combs which encode equations of the form $\tau_{x}+\tau_{y}=C$ for some constant $C$. That allows us to couple $\tau_{x}$ and $\tau_{y}$ to one another such that the two axis-aligned combs cannot be aligned independently from each other.

### v2

@@ Abstract
Given two polygons of complexities $n$ and $m$ respectively, a fundamental problem in shape matching and geometric similarity is to compute their maximum area overlap under translation. For general simple polygons, the best-known algorithm runs in $\mathcal{O}((nm)^{2}\log(nm))$ time [Mount, Silverman, Wu ’96]. In a recent breakthrough that received the SoCG Best Paper Award 2025, Chan and Hair gave a linear-time algorithm for the special case when both polygons are convex. A key challenge in computational geometry is to design improved algorithms for other natural classes of polygons. We address this by presenting an $\mathcal{O}((nm)^{3/2}\log(nm))$-time algorithm for the case when both polygons are orthogonal, probably the most popular class of polygons besides convex and simple ones. This is the first algorithm for polygon overlap on orthogonal polygons that is faster than the almost 30 years old algorithm for general simple polygons.
Complementing our algorithmic contribution, we provide $k$-SUM lower bounds for problems on simple polygons with only orthogonal and diagonal edges. First, we establish that there is no algorithm for polygon overlap with running time $\mathcal{O}(\max(n^{2},nm^{2})^{1-\varepsilon})$, where $m\leq n$, unless the $k$-SUM Hypothesis fails. This matches the running time of our algorithm when $n=m$. We use part of the above construction to also show a lower bound for the polygon containment problem, a popular special case of the overlap problem. Concretely, there is no algorithm for polygon containment with running time $\mathcal{O}(n^{2-\varepsilon})$ under the $3$-SUM Hypothesis, even when the polygon to be contained has $m\in\mathcal{O}(1)$ vertices. Our lower bound shows that polygon containment for these types of polygons (i.e., with diagonal edges) is strictly harder than for orthogonal polygons, and also strengthens the previously known $3$-SUM lower bound for polygon containment of [Barequet, Har-Peled ’01]. Furthermore, our lower bounds show conditional tightness (up to polylogarithmic factors) of the algorithms of [Avnaim, Boissonnat ’89] and [Mount, Silverman, Wu ’96] when $m\in\mathcal{O}(1)$.
@@ 1 Introduction
We consider the problem of computing the maximum overlap of two polygons $P$ and $Q$ with $n$ and $m$ vertices, respectively, i.e., finding a translation $\tau=(\tau_{x},\tau_{y})\in\mathbb{R}^{2}$ such that the area of $P\cap(Q+\tau)$ is maximum. This can be seen as a formalization of matching the shapes $P$ and $Q$, and the maximum overlap is a natural similarity measure: the larger it is relative to the areas of $P$ and $Q$, the more similar they are. Shape matching itself is a core technique in geographic information science, computer aided design, robotics, and medical imaging. Applications include remote sensing, fingerprint matching, molecular shape matching, and image sequence alignment [meng2012fast, choi2007fingerprint, kim2011pubchem3d, 10.1007/3-540-55426-2_53].
The maximum overlap problem for general simple polygons was first studied by Mount, Silverman and Wu [DBLP:journals/cviu/MountSW96]. They show that the overlap function $\tau\mapsto{{\color[rgb]{0,0,0}\mathrm{area}}}(P\cap(Q+\tau))$ is a continuous and piecewise polynomial surface of degree at most two, consisting of ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{2})$ pieces. Furthermore, they give an example of two orthogonal polygons (for arbitrary values of $n$ and $m$) that realize the bound, i.e., the complexity of the overlap function is $\Omega((nm)^{2})$, and hence the bound is tight in the worst-case. They also describe an algorithm to compute the entire function within the same bound ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{2})$ on the running time, which yields an ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{2}\log(nm))$ algorithm for the polygon overlap problem. They posed the open question whether an algorithm with better running time exists. Approximation algorithms for finding the maximum overlap are described in [DBLP:journals/dcg/CheongEH07, DBLP:journals/comgeo/ChengL13, DBLP:journals/algorithmica/Har-PeledR17]. See also the survey [o2017polygons] on polygons from the handbook of discrete and computational geometry for an overview.
@@
Maximizing the overlap has also been studied for convex polyhedra in higher dimensions, with Ahn, Brass, and Shin [ahn2008maximum] providing an algorithm that runs in ${{\color[rgb]{0,0,0}\tilde{\mathcal{O}}}}(n^{d+1-3/d})$ expected time. This was later improved upon by Ahn, Cheng and Reinbacher [ahn2013maximum], who presented an algorithm that for any $\mu>0$ computes a translation for which the area of the overlap is at most the optimum overlap minus $\mu$. The algorithm has running time ${{\color[rgb]{0,0,0}\tilde{\mathcal{O}}}}(n)$ for $d=3$, and ${{\color[rgb]{0,0,0}\tilde{\mathcal{O}}}}(n^{\lfloor d/2\rfloor+1})$ for $d\geq 4$.
Another variant of the problem that has been studied extensively is the minimization of the area of the convex hull of $P\cup(Q+\tau)$. For this problem, Jung, Kang, and Ahn [jung2025minimum] presented an algorithm with deterministic running time $O(n)$ for $d=2$, $O(n^{(d+1)/2})$ for odd $d\geq 3$, and ${{\color[rgb]{0,0,0}\tilde{\mathcal{O}}}}(n^{d/2})$ for even $d\geq 3$.
@@
Convex Containment $\Omega(n+m)$ ${{\color[rgb]{0,0,0}\mathcal{O}}}(n+m)$ [DBLP:conf/compgeom/ChanH25]
MaxOverlap $\Omega(n+m)$ ${{\color[rgb]{0,0,0}\mathcal{O}}}(n+m)$ [DBLP:conf/compgeom/ChanH25]
Orthogonal Containment ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{1-{{\color[rgb]{0,0,0}\mathrm{\varepsilon}}}})$ [KunnemannPolygonPlacement2021] ${{\color[rgb]{0,0,0}\tilde{\mathcal{O}}}}(nm)$ [DBLP:journals/ipl/Barrera96]
MaxOverlap ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{1-{{\color[rgb]{0,0,0}\mathrm{\varepsilon}}}})$ [KunnemannPolygonPlacement2021] ${{\color[rgb]{0,0,0}\tilde{\mathcal{O}}}}((nm)^{3/2})$
@@
Containment ${{\color[rgb]{0,0,0}\mathcal{O}}}(n^{2-{{\color[rgb]{0,0,0}\mathrm{\varepsilon}}}})$ ${{\color[rgb]{0,0,0}\tilde{\mathcal{O}}}}((nm)^{2})$ [DBLP:journals/ita/AvnaimB89]
@@
MaxOverlap ${{\color[rgb]{0,0,0}\mathcal{O}}}(\max(n^{2},nm^{2})^{1-{{\color[rgb]{0,0,0}\mathrm{\varepsilon}}}})$ ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{2})$ [DBLP:journals/ita/AvnaimB89]
@@
In this paper, we consider the maximum overlap problem in ${{\color[rgb]{0,0,0}\mathbb{R}}}^{2}$ for a different well-studied subclass of simple polygons, where both $P$ and $Q$ are orthogonal polygons, i.e., simple polygons where every edge is either horizontal or vertical:
@@
For this problem, we give the first improvement over the nearly three decades old ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{2}\log(nm))$ bound from [DBLP:journals/cviu/MountSW96]. In particular, we prove the following theorem.
@@
The corresponding hardness assumption in fine-grained complexity theory is the $k$-SUM-Hypothesis [abboud2013exact, GAJENTAAN1995165]:
@@
Our reduction can also be modified to obtain a lower bound for the polygon containment problem. While this problem is a special case of the maximum overlap problem, it has attracted more attention than the overlap problem. In the containment problem we want to know whether there is some $\tau\in\mathbb{R}^{2}$ so that $(Q+\tau)\subset P$ and thus ${{\color[rgb]{0,0,0}\mathrm{area}}}(P\cap(Q+\tau))={{\color[rgb]{0,0,0}\mathrm{area}}}(Q)$. This problem was first studied by Chazelle in 1983 [chazelle1983polygon], followed by numerous other results. Of particular interest, we only mention the results when both $P$ and $Q$ are simple or orthogonal polygons and only translations are allowed; a plethora of results exist when rotations are allowed and for more restricted variants such as when one or both of the polygons are convex; for exhaustive literature reviews on the problem, see the recent papers [KunnemannPolygonPlacement2021, DBLP:conf/compgeom/ChanH24]. Avnaim and Boissonnat [DBLP:journals/ita/AvnaimB89] gave an algorithm for solving polygon containment of two simple polygons in ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{2}\log(nm))$ time. Hernández-Barrera [DBLP:journals/ipl/Barrera96] improved the running time to ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{2})$ and solved the case where both $P$ and $Q$ are orthogonal in ${{\color[rgb]{0,0,0}\mathcal{O}}}(nm\log(nm))$ time.
As for known lower bounds, Barequet and Har-Peled [DBLP:journals/ijcga/BarequetHardness01] showed that when $P$ and $Q$ both have $n$ vertices, no algorithm has running time ${{\color[rgb]{0,0,0}\mathcal{O}}}(n^{2-\varepsilon})$ under the $3$-SUM Hypothesis. Künnemann and Nusser [KunnemannPolygonPlacement2021] gave various conditional lower bounds for polygon containment and the more general largest scaled containment problem, which is to find the largest homothet (i.e., scaled copy) of $Q$ that fits in $P$. For polygon containment, they showed that when $P$ and $Q$ are orthogonal and have arbitrary numbers $n$ and $m$ of vertices, then no algorithm has running time ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{1-\varepsilon})$ time under the Orthogonal Vectors Hypothesis. Assuming the $4$-SUM Hypothesis, they showed that for the largest scaled containment problem, no algorithm has running time ${{\color[rgb]{0,0,0}\mathcal{O}}}(n^{2-\varepsilon})$, even when both $P$ and $Q$ are orthogonal and $m=O(1)$.
@@
  • The arrangement that is constructed for the polygon overlap problem in [DBLP:journals/cviu/MountSW96] has worst-case size $\Theta((nm)^{2})$, even for orthogonal polygons. We break this natural barrier with our ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{3/2}\log nm)$ running time.
@@
  • With our improved lower bounds, we show that the decades-old algorithms of [DBLP:journals/ita/AvnaimB89, DBLP:journals/cviu/MountSW96, DBLP:journals/ipl/Barrera96] for polygon overlap and containment are all tight up to polylogarithmic factors for $m\in{{\color[rgb]{0,0,0}\mathcal{O}}}(1)$.
@@ 1.1 Technical Overview
Let $P$ and $Q$ be two orthogonal polygons with $n$ and $m$ vertices respectively. We give the first algorithm improving on the original ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{2})$ algorithm from Mount, Silverman and Wu [DBLP:journals/cviu/MountSW96] for non-convex polygons. That is, for the subclass of orthogonal polygons we give an ${{\color[rgb]{0,0,0}\tilde{\mathcal{O}}}}((nm)^{3/2})$ algorithm.
@@
Similar to the approach presented in [DBLP:journals/cviu/MountSW96], we consider the piecewise polynomial surface of degree at most two describing the function $\tau\mapsto{{\color[rgb]{0,0,0}\mathrm{area}}}(P\cap(Q+\tau))$. We observe that the function splits into ${{\color[rgb]{0,0,0}\mathcal{O}}}(nm)$ rectangles, each with their own bilinear function, and $\tau\mapsto{{\color[rgb]{0,0,0}\mathrm{area}}}(P\cap(Q+\tau))$ can be expressed as the sum of bilinear functions of rectangles containing $\tau$. We may now sweep over these rectangles to compute the translation maximizing the intersection of $P$ and $Q+\tau$ in ${{\color[rgb]{0,0,0}\tilde{\mathcal{O}}}}((nm)^{2})$ time. We improve this sweep-line algorithm with a batching technique which groups events for the sweep-line algorithm into ${{\color[rgb]{0,0,0}\mathcal{O}}}(\sqrt{nm})$ groups of size ${{\color[rgb]{0,0,0}\mathcal{O}}}(\sqrt{nm})$ in both $x$- and $y$-direction resulting in ${{\color[rgb]{0,0,0}\mathcal{O}}}(nm)$ small windows in which we identify the maximizing translation in ${{\color[rgb]{0,0,0}\tilde{\mathcal{O}}}}(\sqrt{nm})$ time as follows: Via algebraic manipulations we rewrite the partial sums of bilinear functions, as they would have been encountered and evaluated by the sweep-line algorithm, into a set $V$ of ${{\color[rgb]{0,0,0}\mathcal{O}}}(\sqrt{nm})$ vectors in ${{\color[rgb]{0,0,0}\mathbb{R}}}^{3}$. Identifying the maximum among the ${{\color[rgb]{0,0,0}\mathcal{O}}}(nm)$ candidate translations in the small window then reduces to ${{\color[rgb]{0,0,0}\mathcal{O}}}(\sqrt{nm})$ extreme-point queries on the set $V$ which, as $V\subset{{\color[rgb]{0,0,0}\mathbb{R}}}^{3}$, can be answered in ${{\color[rgb]{0,0,0}\tilde{\mathcal{O}}}}(\sqrt{nm})$ total time, via the convex hull of $V$. The result is an algorithm with running time in ${{\color[rgb]{0,0,0}\tilde{\mathcal{O}}}}((nm)^{3/2})$.
@@
Comparing our techniques to the algorithm given by Mount, Silverman and Wu [DBLP:journals/cviu/MountSW96], we observe that our approach is similar in spirit in that we consider the piecewise polynomial surface of degree at most two describing the function $\tau\mapsto{{\color[rgb]{0,0,0}\mathrm{area}}}(P\cap(Q+\tau))$. As noted in Section 1, Mount, Silverman and Wu showed that the overlap function can have complexity $\Omega((nm)^{2})$ even for orthogonal polygons, so their method of explicitly computing the polynomial surface is unlikely to yield improved bounds. Unlike their algorithm, however, we manage to circumvent the explicit computation of the polynomial surface represented as the sum of bilinear polynomials defined on rectangular regions via our novel recast to extreme-point queries to a three-dimensional convex hull. With this, we improve the running time of the algorithm given by Mount, Silverman and Wu by a factor of $\sqrt{nm}$.
@@
Comparing our techniques with the techniques of the 3-SUM Hypothesis based lower bound from Barequet and Har-Peled [DBLP:journals/ijcga/BarequetHardness01] and the Orthogonal Vectors Hypothesis based lower bound from Künnemann and Nusser [KunnemannPolygonPlacement2021] for the OrthMaxOverlap problem, we observe the following: Both their techniques essentially consider only one pair of combs, restricting the set $\tau=(\tau_{x},\tau_{y})$ of translations such that $Q+\tau$ is contained in $P$ to one of $nm$ possible $x$-translations. Combining this with a second set of combs restricting the $y$-translations is straight-forward, however, there is no interdependence between $\tau_{x}$ and $\tau_{y}$, as both combs can be aligned independently of one another, giving ${{\color[rgb]{0,0,0}\mathcal{O}}}(nm)$ instead of the desired ${{\color[rgb]{0,0,0}\mathcal{O}}}((nm)^{2})$ translations that ‘need to be checked’. This problem is addressed in the 4-SUM lower bound of [KunnemannPolygonPlacement2021] by considering scalings additionally to translations, to determine the biggest copy of $Q$ in $P$. This additional degree of freedom allows to encode more complexity, however, at the cost of showing a lower bound for a more general problem (hence, a weaker bound). In our work, we address this issue by introducing diagonal combs which encode equations of the form $\tau_{x}+\tau_{y}=C$ for some constant $C$. That allows us to couple $\tau_{x}$ and $\tau_{y}$ to one another such that the two axis-aligned combs cannot be aligned independently from each other. A somewhat similar technique of encoding addition of a 3-SUM problem via diagonal translations was used by Bringmann and Nusser in [DBLP:journals/jocg/BringmannN21].